# Normalized mutual information is a biased measure for classification and community detection

### v1

@@ Abstract
Normalized mutual information is widely used as a similarity measure for evaluating the performance of clustering and classification algorithms. In this paper, we show that results returned by the normalized mutual information are biased for two reasons: first, because they ignore the information content of the contingency table and, second, because their symmetric normalization introduces spurious dependence on algorithm output. We introduce a modified version of the mutual information that remedies both of these shortcomings. As a practical demonstration of the importance of using an unbiased measure, we perform extensive numerical tests on a basket of popular algorithms for network community detection and show that one’s conclusions about which algorithm is best are significantly affected by the biases in the traditional mutual information.
@@ I Introduction
A common task in data analysis is the comparison of two different labelings of the same set of objects. How well do demographics predict political affiliation? How accurately do blood tests predict clinical outcomes? How well do clustering algorithms recover known classes of items? The information theoretic measure known as mutual information is commonly applied to answer questions like these, in which an experimental or computational estimate of some kind is compared against a “ground truth.”
@@
Mutual information CT06 is a measure of how easy it is to describe one labeling of a set of objects if we already know another labeling. Specifically, it measures how much less information it takes (in the Shannon sense) to communicate the first labeling if we know the second versus if we do not. As an example, mutual information is commonly used in network science to evaluate the performance of algorithms for network community detection DDDA05 . One takes a network whose community structure is already known and applies a community detection algorithm to it to infer the communities. Then one uses mutual information to compare the output of the algorithm to the known correct communities. Algorithms that consistently achieve high mutual information scores are considered good. We will use this application as an illustrative example later in the paper.
@@
Mutual information has a number of appealing properties as a tool for comparing labelings. It is invariant under permutations of the labels, so that labelings do not have to be aligned before comparison. It also generalizes gracefully to the case where the number of distinct labels is not the same in the two labelings. On the other hand, the mutual information, in its most common form, also has some significant drawbacks, and two in particular that we highlight in this paper. First, it has a bias towards labelings with too many distinct label values. For instance, a community detection algorithm that routinely divides networks into significantly more groups than are present in the ground truth can nonetheless achieve high mutual information scores. Indeed, if the algorithm simply places every node of the network in a group on its own, the mutual information gives this division a maximum score, regardless of the ground truth. In most cases this would be considered an error. A number of approaches for correcting this flaw have been proposed. One can apply direct penalties for incorrect numbers of groups AP17 or subtract correction terms based on the average value of the mutual information over some ensemble of candidate labelings VEB10 ; GA17 or the statistics of the contingency table NCY20 . For reasons we discuss shortly, we favor the latter approach, which leads to the measure known as the reduced mutual information.
@@
The second drawback of the mutual information arises when the measure is normalized, as it commonly is to improve interpretability. The most popular normalization scheme creates a measure that runs between zero and one by dividing the mutual information by the arithmetic mean of the entropies of the two labelings being compared FJ03 , although one can also normalize by the minimum, maximum, or geometric mean of the entropies. As we demonstrate in this paper, however, these normalizations introduce biases into the results by comparison with the unnormalized measure, because the normalization factor depends on the candidate labeling as well as the ground truth. This effect can be large enough to change the ranking of candidate labelings, and we provide examples of this phenomenon.
@@
In order to avoid this latter bias, while still retaining the interpretability of a normalized mutual information score, we propose normalizing the mutual information by the entropy of the ground-truth labeling alone. This removes the source of bias, at the cost of introducing an asymmetry in the normalization. However, given the asymmetric nature of the classification problems we consider, where one or many candidates are compared against a single ground truth, we argue that this is justified. Furthermore, by contrast with the multitude of possible symmetric normalizations, the measure we propose is the unique unbiased way to normalize the mutual information such that a perfect match receives a score of 1.
@@
Both drawbacks of the standard mutual information can be addressed simultaneously by defining an asymmetrically normalized version of the reduced mutual information, and we advocate for this measure in comparing candidate labelings against a known ground truth. In support of this approach we present an extensive comparison of the performance of this and other variants of the mutual information in network community detection tasks, generating a large number of random test networks with known community structure and a variety of structural parameters then attempting to recover the communities using popular community detection algorithms. Within this framework, we find that conclusions about which algorithms perform best are significantly impacted by the choice of mutual information measure, and specifically that algorithms that find too many communities are erroneously favored by traditional measures.
@@ II Mutual information
Conventionally one uses base-2 logarithms in this context, which gives $H$ units of bits, but we use natural logarithms in this paper, since they are more convenient for our purposes. The only difference is an overall factor of $\ln 2$, which will have no effect on our conclusions. Thus, henceforth, all logarithms should be read as natural logs.
Suppose now that we have a labeling or division of $n$ objects into $q$ groups, which we represent as a vector of $n$ integer elements, each with value in the range $1\ldots q$. We assume there to be a ground-truth labeling, which we denote $g$, and a candidate labeling for comparison $c$, generated for instance by some sort of algorithm. The mutual information $I(c;g)$ between the two is the amount of information that can be saved when transmitting the truth $g$ if the receiver already knows the candidate $c$. We can write this information as the total entropy of $g$ minus the conditional entropy:
@@
It is simple to show that, despite appearances, this definition is symmetric under the interchange of the labelings $c$ and $g$.
We now write down expressions for both entropies in (2), being careful to retain some terms that are often neglected. First consider $H(g)$, which represents the information required to transmit the ground truth on its own. The transmission process has three steps. First, we transmit the number of groups $q_{g}$ in the labeling. The maximum possible value of $q_{g}$ is $n$, so transmitting any particular value requires information $H(q_{g})=\log n$ (see Eq. (1)). Next we transmit a vector $n^{(g)}$ with $q_{g}$ elements $n^{(g)}_{s}$ equal to the number of objects in each group. By definition, the elements of $n^{(g)}$ sum to $n$, and the number of $q_{g}$-element vectors of positive integers that sum to $n$ is $\binom{n-1}{q_{g}-1}$, so the amount of information required to transmit the vector is
$H(n^{(g)}|q_{g})=\log\binom{n-1}{q_{g}-1}.$ (3)
Finally, given $n^{(g)}$, there are $n!/\prod_{s}n_{s}^{(g)}!$ labelings $g$ that have the correct group sizes, so the amount of information needed to identify a particular labeling uniquely is
@@
The total cost to transmit $g$ is equal to the sum of the individual costs of the three steps, thus:
$H(g)$ $=H(q_{g})+H(n^{(g)}|q_{g})+H(g|n^{(g)})$
$=\log n+\log\binom{n-1}{q_{g}-1}+\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}.$ (5)
@@
This three-part encoding scheme is not the only one that could be applied to this problem, but it is an efficient one in the common case of a relatively small number of groups $q_{g}\ll n$ with potentially unequal sizes, and it is the one upon which the conventional definition of mutual information is based, which is our focus here.
@@
To calculate the second term in Eq. (2) we use a similar multipart scheme, but one that now takes advantage of the receiver’s knowledge of $c$ when communicating $g$. We first communicate $q_{g}$ and $n^{(g)}$ as before, at the same information cost of $H(q_{g})+H(n^{(g)}|q_{g})=\log n+\log\binom{n-1}{q_{g}-1}$. Then we communicate a contingency table $n^{(cg)}$, a matrix with elements $n_{rs}^{(cg)}$ equal to the number of objects that simultaneously belong to group $r$ in the candidate labeling $c$ and group $s$ in the ground truth $g$. Figure 1 shows an example of a contingency table for two labelings of the nodes in a small network.
@@
Figure 1: An example of a contingency table for two colorings of the same network, one with three colors and one with two. The entries in the $3\times 2$ contingency table $n^{(cg)}$ count the number of nodes that have each combination of colors. The row and column sums $n^{(c)}$ and $n^{(g)}$ of the contingency table count the number of nodes with each color in the two colorings. Note that, although we illustrate the contingency table with an application to a network, the table itself is independent of the network structure.
The contingency table consists of non-negative integer elements and its row and column sums give the sizes of the groups in $c$ and $g$ respectively:
$\sum_{s=1}^{q_{g}}n_{rs}^{(cg)}=n_{r}^{(c)},\qquad\sum_{r=1}^{q_{c}}n_{rs}^{(cg)}=n_{s}^{(g)}.$ (6)
Since the receiver already knows the values of $n^{(c)}$ and $n^{(g)}$ (the former because they know $c$ and the latter because we just transmitted it), only contingency tables with these row and column sums need be considered. The information cost to transmit the contingency table is thus equal to $\log\Omega(n^{(c)},n^{(g)})$, where $\Omega(n^{(c)},n^{(g)})$ is the number of possible tables with the required row and column sums. There is no known general expression for this number, but approximations exist that are good enough for practical purposes DE85 ; BH10 ; JKN22 .
@@
Finally, having transmitted the contingency table, it remains only to transmit the ground-truth labeling itself, where we need consider only those labelings consistent with the contingency table, given the known candidate labeling $c$. The number of such labelings is $\prod_{r}n_{r}^{(c)}!/\prod_{rs}n_{rs}^{(cg)}!$, so the information needed to uniquely identify one of them is
$H(g|n^{(cg)},c)=\log\frac{\prod_{r}n_{r}^{(c)}!}{\prod_{rs}n_{rs}^{(cg)}!}.$ (7)
Putting everything together, the total conditional information is then
$H(g|c)$ $=H(q_{g})+H(n^{(g)}|q_{g})$
$\qquad{}+H(n^{(cg)}|n^{(c)},n^{(g)})+H(g|n^{(cg)},c)$
$=\log n+\log\binom{n-1}{q_{g}-1}$
$\qquad{}+\log\Omega(n^{(c)},n^{(g)})+\log\frac{\prod_{r}n_{r}^{(c)}!}{\prod_{rs}n_{rs}^{(cg)}!}.$ (8)
@@
In typical applications the number of labelings compatible with the contingency table is much smaller than the total number of labelings, and hence this encoding scheme substantially reduces the amount of information needed to transmit the ground truth. This is not the only encoding possible, but it is an efficient one in practice and it is the one upon which the conventional mutual information is based.
Substituting our expressions (5) and (8) for $H(g)$ and $H(g|c)$ into Eq. (2), we now get the mutual information
@@
$=\log\frac{n!\prod_{rs}n_{rs}^{(cg)}!}{\prod_{r}n_{r}^{(c)}!\prod_{s}n_{s}^{(g)}!}-\log\Omega(n^{(c)},n^{(g)}).$ (9)
Note that the terms $\log n$ and $\log\binom{n-1}{q_{g}-1}$ have cancelled and so play no further role. In most treatments the term $\log\Omega(n^{(c)},n^{(g)})$ for the information content of the contingency table is also dropped, since it is subdominant and often small. When this term is retained, the quantity in (9) is known as the reduced mutual information, and there are good reasons to use this full expression in certain circumstances, a point that we investigate in Section II.1. If we follow convention, however, and neglect the contingency table term, then we arrive at the traditional mutual information, which we denote $I_{0}$:
$I_{0}(c;g)=\log\frac{n!\prod_{rs}n_{rs}^{(cg)}!}{\prod_{r}n_{r}^{(c)}!\prod_{s}n_{s}^{(g)}!}.$ (10)
@@
Note that these differ from the entropy of Eq. (5), which includes the subdominant terms. Trivially (11) and (12) provide upper bounds on the mutual information, while zero is a lower bound, so
@@
$I_{0}(c;g)=\sum_{rs}p_{rs}^{(cg)}\log\frac{p_{rs}^{(cg)}}{p_{r}^{(c)}p_{s}^{(g)}}+\mathrm{O}\left(\frac{\log n}{n}\right),$ (15)
@@
$p_{rs}^{(cg)}={n_{rs}^{(cg)}\over n},\qquad p_{r}^{(c)}={n_{r}^{(c)}\over n},\qquad p_{s}^{(g)}={n_{s}^{(g)}\over n},$ (16)
@@
II.1 Variants of the mutual information
Equation (10) defines the standard mutual information. In this paper we explore variant definitions that can perform better at the classification tasks we consider. A number of variants have been proposed, of which the simplest in the context of the current presentation is the reduced mutual information $I(c;g)$ of Eq. (9), which is simply the value of the mutual information when one does not neglect the term in $\log\Omega(n^{(c)},n^{(g)})$ that represents the cost of transmitting the contingency table:
$I(c;g)=I_{0}(c;g)-\log\Omega(n^{(c)},n^{(g)}).$ (17)
The moniker “reduced” derives from the fact that the contingency table term is always negative and so reduces the value of the mutual information relative to the conventional definition $I_{0}(c;g)$, but we emphasize that functionally we are simply retaining a term that is usually dropped.
As mentioned previously, there is no general closed-form expression for the number $\Omega(n^{(c)},n^{(g)})$ of contingency tables with given row and column sums, and its numerical computation is #P-hard in general DKM97 and hence intractable for all but the smallest of examples. In practice, therefore, the number must be approximated. In this paper we make use of the “effective columns” approximation of JKN22 , which has good performance over a wide range of situations and a simple closed-form expression:
$\Omega(n^{(c)},n^{(g)})\simeq$
$\binom{n+q_{c}\alpha-1}{q_{c}\alpha-1}^{-1}\prod_{r=1}^{q_{c}}\!\binom{n_{r}^{(c)}+\alpha-1}{\alpha-1}\prod_{s=1}^{q_{g}}\!\binom{n_{s}^{(g)}+q_{c}-1}{q_{c}-1},$ (18)
where
$\alpha=\frac{n^{2}-n+\bigl{[}n^{2}-\sum_{s}(n_{s}^{(g)})^{2}\bigr{]}/q_{c}}{\sum_{s}(n_{s}^{(g)})^{2}-n}.$ (19)
This estimate differs from the one originally used for the reduced mutual information in Newman06b , but we favor it here since it performs better in certain regimes.
To understand the importance of the contingency table term in the mutual information, let us explore how things can go wrong if we neglect it. To take an extreme example, suppose an algorithm simply places every object in a group of its own, which corresponds to the candidate labeling $c=(1,\ldots,n)$. No matter what the ground truth labeling is, this choice of $c$ clearly contains no information about it whatsoever, so we might expect the mutual information to be zero. The contingency table in this case is
$n_{rs}^{(cg)}=\biggl{\{}\begin{array}[]{ll}1&\quad\mbox{if $g_{r}=s$,}\\ 0&\quad\mbox{otherwise,}\\ \end{array}$ (20)

### v2

@@ Abstract
Normalized mutual information is widely used as a similarity measure for evaluating the performance of clustering and classification algorithms. In this paper, we argue that results returned by the normalized mutual information are biased for two reasons: first, because they ignore the information content of the contingency table and, second, because their symmetric normalization introduces spurious dependence on algorithm output. We introduce a modified version of the mutual information that remedies both of these shortcomings. As a practical demonstration of the importance of using an unbiased measure, we perform extensive numerical tests on a basket of popular algorithms for network community detection and show that one’s conclusions about which algorithm is best are significantly affected by the biases in the traditional mutual information.
@@ I Introduction
A common task in data analysis is the comparison of two different labelings of the same set of objects. How well do demographics predict political affiliation? How accurately do blood tests predict clinical outcomes? The information theoretic measure known as mutual information is commonly applied to answer questions like these, in which an experimental or computational estimate of some kind is compared against a “ground truth.”
@@
Mutual information [1] is a measure of how efficiently we can describe one labeling of the set of objects if we already know another labeling. Specifically, it measures how much less information it takes to communicate the first labeling if we know the second versus if we do not. As an example, mutual information is commonly used in network science to evaluate the performance of algorithms for network community detection [2]. One takes a network whose community structure is already known and applies a community detection algorithm to it to infer the communities. Then one uses mutual information to compare the output of the algorithm to the known correct communities. Algorithms that consistently achieve high mutual information scores are considered good. We will use this application as an illustrative example later in the paper.
@@
Mutual information has a number of appealing properties as a tool for comparing labelings. It is invariant under permutations of the labels, so that, for example, two divisions of a network into communities that differ only in how the communities are labeled will be correctly identified as being the same. It also returns sensible results in cases where the number of distinct label values is not the same in the two labelings. However, the standard mutual information measure also has some significant shortcomings, and two in particular that we highlight in this paper. First, it has a bias towards labelings with too many distinct label values. For instance, a community detection algorithm that routinely divides networks into significantly more groups than are present in the ground truth can nonetheless achieve high mutual information scores. A number of approaches for correcting this flaw have been proposed. One can apply direct penalties for incorrect numbers of groups [3] or subtract correction terms based on the average value of the mutual information over some ensemble of candidate labelings [4, 5] or on the statistics of the contingency table [6, 7]. For reasons we discuss shortly, we favor the latter approach, which leads to the measure known as the reduced mutual information.
@@
The second drawback of the mutual information arises when the measure is normalized, as it commonly is to improve interpretability. The most popular normalization scheme creates a measure that runs between zero and one by dividing the mutual information by the arithmetic mean of the entropies of the two labelings being compared [8], although one can also normalize by the minimum, maximum, or geometric mean of the entropies. As we demonstrate in this paper, however, these normalizations introduce biases into the results by comparison with the unnormalized measure, because the normalization factor depends on the candidate labeling as well as the ground truth. This effect can be large enough to change scientific conclusions, and we provide examples of this phenomenon.
@@
In order to avoid this latter bias, while still retaining the interpretability of a normalized mutual information measure, in this paper we propose normalizing by the entropy of the ground-truth labeling alone. This removes the source of bias but introduces an asymmetry in the normalization. At first sight this asymmetry might seem undesirable, and previous authors have gone to some lengths to avoid it. Here, however, we argue that it is not only justified but even desirable, for several reasons. First, many of the classification problems we consider are unaffected by the asymmetry, since they involve the comparison of one or more candidate labelings against a single, unchanging ground truth. Moreover, by contrast with the multitude of possible symmetric normalizations, the asymmetric measure we propose is the unique unbiased way to normalize the mutual information such that a perfect match receives a score of 1.
But more broadly, one can argue that an asymmetric measure can be more informative than the conventional symmetric one. Consider for instance the common situation where one labeling is a fine-grained version of the other. We might for example label individuals by the country they live in on the one hand and by the town or city on the other. Then the detailed labeling tells us everything there is to know about the coarse-grained one, but the reverse is not true: telling you the city fixes the country, but not vice versa. Thus one could argue that the mutual information between the two should be asymmetric even before normalization—labeling A contains more information about labeling B than B does about A—and this type of asymmetry will be a key feature of the measures we study.
@@
Both drawbacks of the standard mutual information—bias towards too many groups and dependence of the normalization on the candidate labeling—can be addressed simultaneously by using an asymmetric normalized reduced mutual information as defined in this paper. In support of this approach we present an extensive comparison of the performance of this and other variants of the mutual information in network community detection tasks, generating a large number of random test networks with known community structure and a variety of structural parameters, and then attempting to recover the communities using popular community detection algorithms. Within this framework, we find that conclusions about which algorithms perform best are significantly impacted by the choice of mutual information measure, and specifically that traditional measures erroneously favor algorithms that find too many communities, but our proposed measure does not. Code implementing our measure may be found at https://github.com/maxjerdee/reduced_mutual_information.
@@ II Mutual information
Conventionally one uses base-2 logarithms in this context, which gives $H$ units of bits, but we use natural logarithms in this paper, since they are more convenient for our purposes. The only difference is an overall factor of $\ln 2$, which will have no effect on our conclusions. Henceforth, all logarithms should be read as natural logs.
Suppose that we have a labeling or division of $n$ objects into $q$ groups, which we represent as a vector of $n$ integer elements, each with value in the range $1\ldots q$. We assume there to be a ground-truth labeling, which we denote $g$, and a candidate labeling for comparison $c$, generated for instance by some sort of algorithm. The mutual information $I(c;g)$ between the two is the amount of information that can be saved when transmitting the truth $g$ if the receiver already knows the candidate $c$. We can write this information as the total entropy of $g$ minus the conditional entropy of $g$ given $c$:
@@
Loosely, we say that $I(c;g)$ measures “how much $c$ tells us about $g$.”
We now write down expressions for both entropies in (2), but there are some subtleties involved in doing so, because the expressions depend on how the transmitted information is encoded and also because achieving correct results requires us to carefully retain some terms that are often neglected. Our presentation follows that of [7]. First consider $H(g)$, which represents the information required to transmit the ground truth on its own. The transmission process has three steps. We first transmit the number of groups $q_{g}$ in the labeling, followed by a vector $n^{(g)}$ with $q_{g}$ elements $n^{(g)}_{s}$ equal to the number of objects in each group. Third, we transmit our final labeling $g$. If we apply Eq. (1) to calculate the cost of the first step, since the maximum possible value of $q_{g}$ is $n$, transmitting any particular value requires information $H(q_{g})=\log n$. Likewise for the third step, given $n^{(g)}$, there are $n!/\prod_{s}n_{s}^{(g)}!$ labelings $g$ that have the correct group sizes, so the amount of information needed to identify a particular labeling uniquely is
@@
The middle step, transmission of the group sizes, is more delicate: here some optimization is possible that allows us to achieve a better measure of information content. Generally when transmitting a sequence of $n$ values with unequal probabilities such that value $r$ occurs $n_{r}$ times, the information cost is given by the generalization of Eq. (3) thus:
$\log{n!\over\prod_{r}n_{r}!}$ $\simeq n\log n-n-\sum_{r}(n_{r}\log n_{r}-n_{r})$
$=-\sum_{r}n_{r}\log p_{r},$ (4)
where $p_{r}=n_{r}/n$ is the probability of value $r$ and we have approximated the factorials using Stirling’s formula. Equation 4 tells us that the information cost to transmit the value $r$ is simply $-\log p_{r}$. If the $N$ possible values are equally likely to appear and $p_{r}=1/N$, we recover the information cost of Eq. (1), but if they are not we can take advantage of that fact to create more efficient encodings.
The transmission of $n^{(g)}$ benefits from such an efficient encoding. We do not in general know the distribution of the elements of $n^{(g)}$, but we know they that may be—and often are—heterogeneous. As shown in [7], an effective approach is to assume a Dirichlet-multinomial distribution with concentration parameter $\alpha_{g}>0$
$P(n^{(g)}|q_{g},\alpha_{g})=$ (5)
$\qquad\binom{n+q_{g}\alpha_{g}-1}{q_{g}\alpha_{g}-1}^{-1}\prod_{r=1}^{q_{g}}\binom{n_{r}^{(g)}+\alpha_{g}-1}{\alpha_{g}-1},$
where for non-integer $\alpha_{g}$ we generalize the binomial coefficient in the obvious way:
$\Bigl{(}{n\atop k}\Bigr{)}={\frac{\Gamma(n+1)}{\Gamma(k+1)\Gamma(n-k+1)}},$ (6)
with $\Gamma(x)$ being the standard gamma function. If $\alpha_{g}=1$, the Dirichlet-multinomial distribution is uniform over all $q_{g}$-vectors $n^{(g)}$ of non-negative integers that sum to $n$. If $10<\alpha_{g}<$ 1 the distribution favors vectors with more extreme elements, while for $\alpha_{g}>1$ it favors vectors with more homogeneous elements. With this choice the resulting information cost to transmit $n^{(g)}$ is
$H($ $n^{(g)}|q_{g},\alpha_{g})=-\log P(n^{(g)}|q_{g},\alpha_{g})$ (7)
$=\log\binom{n+q_{g}\alpha_{g}-1}{q_{g}\alpha_{g}-1}-\sum_{r=1}^{q_{g}}\log\binom{n_{r}^{(g)}+\alpha_{g}-1}{\alpha_{g}-1}.$
The optimal encoding for transmitting $n^{(g)}$ within the Dirichlet-multinomial family is given by the minimum of this expression with respect to $\alpha_{g}$, which is also equivalent to simply maximizing $P(n^{(g)}|\alpha_{g})$, i.e., to finding the maximum-likelihood value of $\alpha_{g}$. Note that we omit the cost of transmitting the parameter $\alpha_{g}$ itself since it will ultimately cancel out of the expression for the mutual information as described in [7].
@@
The total cost to transmit $g$ is now equal to the sum of the individual costs of the three steps, thus:
$H(g)$ $=H(q_{g})+H(n^{(g)}|q_{g},\alpha_{g})+H(g|n^{(g)})$
$=\log n+\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}$ (8)
$+\log\binom{n+q_{g}\alpha_{g}-1}{q_{g}\alpha_{g}-1}-\sum_{r=1}^{q_{g}}\log\binom{n_{r}^{(g)}+\alpha_{g}-1}{\alpha_{g}-1}$
@@
This three-part encoding scheme is not the only one that can be applied to this problem, but it is an efficient one in the common case of a relatively small number of groups $q_{g}\ll n$ with potentially unequal sizes, and it reduces to the conventional expression for $H(g)$ when subleading terms are neglected. It is an important part of our argument here, however, that the subleading terms be retained and carried through to the calculation of the mutual information.
@@
For the calculation of $H(g|c)$, the second term in Eq. (2), we use a similar three-part encoding, but one that takes advantage of the receiver’s knowledge of $c$ when communicating $g$. We first communicate $q_{g}$ as before, at the same information cost of $H(q_{g})=\log n$. In place of the vector $n^{(g)}$ we then communicate a contingency table $n^{(gc)}$, a matrix with elements $n_{rs}^{(gc)}$ equal to the number of objects that simultaneously belong to group $r$ in the candidate labeling $c$ and group $s$ in the ground truth $g$. Figure 1 shows an example of a contingency table for two labelings of the nodes in a small network.
@@
Figure 1: An example of a contingency table for two colorings of the same network, one with three colors and one with two. The entries in the $3\times 2$ contingency table $n^{(gc)}$ count the number of nodes that have each combination of colors. The row and column sums $n^{(g)}$ and $n^{(c)}$ of the contingency table count the number of nodes with each color in the two colorings. Note that, although we illustrate the contingency table with an application to a network, the table itself is independent of the network structure.
Our process for transmitting the contingency table is analogous to the one we use for transmitting $n^{(g)}$. Since the labeling $c$ is known, the column sums $n^{(c)}$ of the contingency table are known as well, and we assume each column to be drawn from a Dirichlet-multinomial distribution with the appropriate sum. We use the same concentration parameter $\alpha_{g|c}$ for every column, but the columns are otherwise independent. If we denote column $s$ by $n_{\cdot s}^{(gc)}$, then the information cost to transmit the entire contingency table is
$H($ $n^{(gc)}|n^{(c)},q_{g},\alpha_{g|c})=\sum_{s=1}^{q_{c}}H(n_{\cdot s}^{(gc)}|n_{s}^{(c)},q_{g},\alpha_{g|c})$
$=\sum_{s=1}^{q_{c}}\biggl{[}\log\binom{n_{s}^{(c)}+q_{g}\alpha_{g|c}-1}{q_{g}\alpha_{g|c}-1}$
$\hskip 60.00009pt-\sum_{r=1}^{q_{g}}\log\binom{n_{rs}^{(gc)}+\alpha_{g|c}-1}{\alpha_{g|c}-1}\biggr{]}.$ (9)
This formulation allows for very efficient transmission of the contingency table under commonly occurring conditions. Consider for instance the special (but not implausible) case where $g=c$, so that $n^{(gc)}$ is diagonal. In this case the optimal concentration parameter is $\alpha_{g|c}=0$, which restricts the possible contingency tables to those for which each column $n_{\cdot s}^{(gc)}$ has exactly one nonzero entry. Given that we know the column sums, this leaves only the positions of the nonzero entries within each column to be transmitted, which results in near-optimal performance, a decisive improvement over previous reductions of the mutual information [7]. Most practical applications will not have $c=g$ precisely, but the two labelings are often similar and similarly impressive improvements in performance can be realized in such cases as well.
@@
Finally, having transmitted the contingency table, it remains only to transmit the ground-truth labeling itself, where we need consider only those labelings consistent with the contingency table, given the known candidate labeling $c$. The number of such labelings is $\prod_{r}n_{r}^{(c)}!/\prod_{rs}n_{rs}^{(gc)}!$, so the information needed to uniquely identify one of them is
$H(g|n^{(gc)},c)=\log\frac{\prod_{r}n_{r}^{(c)}!}{\prod_{rs}n_{rs}^{(gc)}!}.$ (10)
@@
In typical applications the number of labelings compatible with the contingency table is much smaller than the total number of labelings, and hence this encoding substantially reduces the amount of information needed to transmit the ground truth. It is not the only encoding possible, but it is an efficient one in practice and it is the one upon which the conventional mutual information is based.
Collecting all our terms, the total conditional information can now be written
$H(g|c)$ $=H(q_{g})+H(n^{(gc)}|n^{(c)},q_{g},\alpha_{g|c})+H(g|n^{(cg)},c).$ (11)
And substituting (8) and (11) into Eq. (2), we get a complete expression for the mutual information
@@
$=\log\frac{n!\prod_{rs}n_{rs}^{(gc)}!}{\prod_{r}n_{r}^{(c)}!\prod_{s}n_{s}^{(g)}!}+\log\binom{n+q_{g}\alpha_{g}-1}{q_{g}\alpha_{g}-1}$
$\qquad-\sum_{r=1}^{q_{g}}\log\binom{n_{r}^{(g)}+\alpha_{g}-1}{\alpha_{g}-1}$
$\qquad-\sum_{s=1}^{q_{c}}\log\binom{n_{s}^{(c)}+q_{g}\alpha_{g|c}-1}{q_{g}\alpha_{g|c}-1}$
$\qquad+\sum_{r=1}^{q_{g}}\sum_{s=1}^{q_{c}}\log\binom{n_{rs}^{(gc)}+\alpha_{g|c}-1}{\alpha_{g|c}+1}.$ (12)
In most treatments, only the first of these terms is retained, since the others are subdominant and often small [6, 7]. When all the terms are retained, however, the quantity (12) is known as the reduced mutual information and there are good reasons to use this full expression in certain circumstances, a point that we investigate in Section II.1.
If we follow convention, however, and keep only the leading term, then we arrive at the traditional mutual information, which we denote $I_{0}$:
$I_{0}(c;g)=\log\frac{n!\prod_{rs}n_{rs}^{(gc)}!}{\prod_{r}n_{r}^{(c)}!\prod_{s}n_{s}^{(g)}!}.$ (13)
@@
Note that these differ from the entropy of Eq. (8), which includes subdominant terms. Trivially (14) and (15) provide upper bounds on the mutual information, while zero is a lower bound, so
@@
$I_{0}(c;g)=n\sum_{rs}p_{rs}^{(gc)}\log\frac{p_{rs}^{(gc)}}{p_{r}^{(c)}p_{s}^{(g)}}+\mathrm{O}\left(\log n\right),$ (18)
@@
$p_{rs}^{(gc)}={n_{rs}^{(gc)}\over n},\qquad p_{r}^{(c)}={n_{r}^{(c)}\over n},\qquad p_{s}^{(g)}={n_{s}^{(g)}\over n},$ (19)
@@
II.1 Reduced mutual information
Equation (13) defines the standard mutual information. In this paper we explore variant definitions that can perform better at the clustering tasks we consider. A number of variants have been proposed, of which the simplest in the context of the current presentation is the reduced mutual information $I(c;g)$ of Eq. (12), which is simply the value of the mutual information when one does not neglect the subleading information cost of transmitting the group sizes and the contingency table.
To understand the importance of the subleading terms let us explore how things can go wrong if we neglect them. To take an extreme example, suppose an algorithm simply places every object in a group of its own, which corresponds to the candidate labeling $c=(1,\ldots,n)$. No matter what the ground truth is, this choice of $c$ clearly contains no information about it whatsoever, so we might expect the mutual information to be zero. This, however, is not what we find. The contingency table in this case is
$n_{rs}^{(gc)}=\biggl{\{}\begin{array}[]{ll}1&\quad\mbox{if $g_{r}=s$,}\\ 0&\quad\mbox{otherwise,}\\ \end{array}$ (20)